Calculate the number of photons emitted per second by 25 W source of monochromatic light of wavelength 600nm.

Grade:12

1 Answers

Arun
5 years ago
6000Å = 6000^-10m = 6.0^-7m 

From Planck's equation .. each photon has energy, E = hc/λ (J) 

If N photons are emitted per sec ..  
Total emission power, P = N(s⁻¹) x E(J) = NE (J/s) = 25 W 

25 = NE = N x hc/λ  
N = 25λ / hc .. 25 x 6.0^-7 / (6.625^-34 x 3.0^8) .. .. ►N = 7.55^19 photons per sec